The acceleration of an electron in the first orbit of
Hydrogen atom is

Grade:12

1 Answers

Arun
4 years ago

Bohr quantisation principle -

mvr=\frac{nh}{2\pi } \\2\pi r= n\lambda

- wherein

Angular momentum of an electron in  stationary orbit is quantised.

 

 Acceleration =\frac{v^{2}}{r}

\because\ \;mvr=\frac{nh}{2\pi}\ \; \;\Rightarrow\ \;v=\frac{h}{2\pi mr}(n=1)

\therefore\ \; a=\left(\frac{h}{2\pi mr} \right )^{2}.\frac{1}{r}=\frac{h^{2}}{4\pi^{2}m^{2}r^{3}}